OpenStudy (anonymous):

Consider the two vectors A(with arrow above)=5i-3j and B(with arrow above)=-i-2j 1. Calculate A (with arrow above) + B (with arrow above) 2. Calculate A (with arrow above) - B (with arrow above) 3. Calculate absolute value of A (with arrow above) + B (with arrow above) 4. Calculate absolute value of A (with arrow above) - B (with arrow above) 5. Calculate the directions of A (with arrow above) + B (with arrow above) and A (with arrow above) - B (with arrow above). (answers are counterclockwise from the +xaxis)

OpenStudy (anonymous):

i know numbers 1-4..i just dont know question 5..please help!

OpenStudy (anonymous):

COUNTER CLOCKWISE MEANS ANTI-CLOCKWISE

OpenStudy (anonymous):

i know that..but i dont know how to do that direction

OpenStudy (anonymous):

i just dont get number 5 =( i got everything else right

OpenStudy (anonymous):

take the inverse tangent to get the angle

OpenStudy (anonymous):

what is A + B?

OpenStudy (anonymous):

I MEAN -6I+J

OpenStudy (anonymous):

really?

OpenStudy (anonymous):

hold on. unless i am very confused \[5i-3j+(-i-2j)=4i-5j\] correct me if i am wrong

OpenStudy (anonymous):

you are right

OpenStudy (anonymous):

then a-b is 6i-j then for the absolute values i got \[\sqrt{41}\] and \[\sqrt{35}\]

OpenStudy (anonymous):



OpenStudy (anonymous):

so to do the directions...i have to do what exactly?

OpenStudy (anonymous):

you want that angle to find it you see that \[\tan(\theta)=-\frac{5}{4}\] so \[\theta =\tan^{-1}(-\frac{5}{4})\]

OpenStudy (anonymous):

that will give you a negative number. if that annoys you subtract it form 360 to get the positive coterminal angle

OpenStudy (anonymous):

idk what it wants...but thats exactly what it says..i copied the exact question

OpenStudy (anonymous):

assuming you are working in degrees. yes?

OpenStudy (anonymous):

you need a calculator. make sure it is in degree mode and hit second - tan - -5/4

OpenStudy (anonymous):

so for a+b i got -51.34..so to make it positive i will add 360..right?

OpenStudy (anonymous):

yes

OpenStudy (anonymous):

i got 308.66

OpenStudy (anonymous):

i get -51.34 so if you subtract from 360 you get what you wrote

OpenStudy (anonymous):

so the next one will be arc tan of -1/6??

OpenStudy (anonymous):

rather "add to 360"

OpenStudy (anonymous):

yea when i added i got 308.66

OpenStudy (anonymous):

yes the next one would be arctan(-1/6)

OpenStudy (anonymous):

+360

OpenStudy (anonymous):

350.54?

OpenStudy (anonymous):

so it says i got absolute value of a-b wrong

OpenStudy (anonymous):

but everything else is right lol

OpenStudy (anonymous):

really? why isn't it \[\sqrt{37}\]?

OpenStudy (anonymous):

i got \[\sqrt{35}\]

OpenStudy (anonymous):

i added wrong =( lol sorry ill try 37

OpenStudy (anonymous):

yeah it is 36 + 1 not 36 - 1

OpenStudy (anonymous):

thanks!! =) got it all right!..now onto the next question! thanks again!

OpenStudy (amistre64):

A =<5,-3> and B = <-1,-2> 1. Calculate A + B A = <5,-3> +B = <-1,-2> ------------ <4,-5> 2. Calculate A - B A = <5,-3> -B = <-1,-2> ------------ <6,-1> 3. Calculate |A + B| sqrt(4^2 + 5^2) = sqrt(16+25) = sqrt (41) 4. Calculate |A - B| sqrt(6^2 + 1^2) = sqrt(36+1) = sqrt (37) 5. Calculate the directions of A+B and A-B. A+B = <4,-5> pi + tan-1(-5/4) A-B = <6,-1> 2pi - tan-1(-1/6)

OpenStudy (amistre64):

just for practice :)